A jeweler has five rings, each weighing 18 g, made of an alloy of 10% silver adn 90% gold. He decides to melt down the rings and add enough silver to reduce the gold content to 75%. How much silver should he add?

Answer by Paul(988)

10(18)+90(18)=75(x+18)
180+1620=75x+1350
1800=75x+1350
75x=450
x=6

Hence he should add 6g of silver.
Paul.